# TALEC: Teach Your LLM to Evaluate in Specific Domain with In-house Criteria by Criteria Division and Zero-shot Plus Few-shot

### v1

@@ Appendix B Experimental Results
Method | task | label | Spearman/Pearson/Kendall | Acc | F1/Precision/Recall | |||
eval | test | eval | test | eval | test | |||
Arbitrariness | Sentiment Analysis | Not Meeting the Requirements | 0.9551 0.9523 0.9484 | 0.9515 0.943 0.9268 | 0.9123 | 0.8829 | 0.75 / 0.625 / 0.9375 | 0.6286 / 0.4783 / 0.9167 |
Incorrect Answer/Unrelated Matching Results | 0.9737 | 0.9279 | 0.9684 / 1.0 / 0.9388 | 0.92 / 0.9388 / 0.902 | ||||
Refusal to Answer | 0.9737 | 0.964 | 0.6667 / 0.75 / 0.6 | 0.5 / 0.3333 / 1.0 | ||||
Untranslated Text | 0.9825 | 0.982 | 0.875 / 0.7778 / 1.0 | 0.875 / 0.7778 / 1.0 | ||||
Confusing Answers | 0.9912 | 0.982 | 0.8571 / 0.75 / 1.0 | 0.75 / 0.6 / 1.0 | ||||
Stiffness | 0.9649 | 0.9279 | 0.8182 / 0.6923 / 1.0 | 0.6667 / 0.5 / 1.0 | ||||
Repetitive Expression | 0.9649 | 0.955 | 0.7143 / 1.0 / 0.5556 | 0.5455 / 1.0 / 0.375 | ||||
Subject Imprecision | 0.8947 | 0.7838 | 0.7857 / 0.7333 / 0.8462 | 0.3333 / 0.2069 / 0.8571 | ||||
Knowledge QA | Incorrect Answer/Unrelated Matching Results | 0.5242 0.5345 0.4974 | 0.5459 0.5436 0.518 | 0.8085 | 0.7591 | 0.4906 / 0.3333 / 0.9286 | 0.2979 / 0.1944 / 0.6364 | |
Refusal to Answer | 1 | 0.9927 | 1.0/1.0/1.0 | 0.8571 / 0.75 / 1.0 | ||||
Untranslated Text | 0.9858 | 0.9927 | 0.75 / 0.75 / 0.75 | 0.8571 / 1.0 / 0.75 | ||||
Confusing Answers | 0.9858 | 1 | 0.75 / 0.6 / 1.0 | 1.0/1.0/1.0 | ||||
Incomplete Answers | 0.7589 | 0.6934 | 0.6136 / 0.7941 / 0.5 | 0.5625 / 0.871 / 0.4154 | ||||
Stiffness | 0.9716 | 1 | 0.7143 / 0.8333 / 0.625 | 1.0/1.0/1.0 | ||||
Repetitive Expression | 0.9858 | 0.9781 | 0.8571 / 0.8571 / 0.8571 | 0.7273 / 0.8 / 0.6667 | ||||
Subject Imprecision | 0.9645 | 0.9416 | 0.7619 / 0.8889 / 0.6667 | 0.3333 / 0.4 / 0.2857 | ||||
Search QA | Incorrect Answer/Unrelated Matching Results | 0.7946 0.7901 0.7852 | 0.814 0.818 0.791 | 0.9449 | 0.9587 | 0.7586 / 0.7857 / 0.7333 | 0.8387 / 0.9286 / 0.7647 | |
External Links or Diversions | 0.9921 | 0.9752 | 0.9333 / 1.0 / 0.875 | 0.7273 / 0.6667 / 0.8 | ||||
Refusal to Answer | 0.9685 | 1 | 0.7143 / 0.5556 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Incomplete Answers | 0.8661 | 0.9174 | 0.32 / 0.3636 / 0.2857 | 0.5455 / 0.8571 / 0.4 | ||||
Stiffness | 0.9606 | 0.9752 | 0.4444 / 0.2857 / 1.0 | 0.6667 / 0.75 / 0.6 | ||||
Repetitive Expression | 0.9528 | 0.9421 | 0.4 / 0.2857 / 0.6667 | 0.3636 / 0.2222 / 1.0 | ||||
Subject Imprecision | 0.9528 | 0.9174 | 0.5714 / 0.4 / 1.0 | 0.5455 / 0.6 / 0.5 | ||||
Title Generation | Not Meeting the Requirements | 0.535 0.5629 0.5153 | 0.5729 0.593 0.5483 | 0.617 | 0.6147 | 0.6457 / 0.7387 / 0.5734 | 0.6426 / 0.7477 / 0.5634 | |
Incorrect Answer/Unrelated Matching Results | 0.9064 | 0.8745 | 0.8036 / 0.7377 / 0.8824 | 0.7752 / 0.7937 / 0.7576 | ||||
External Links or Diversion | 1 | 1 | 1.0/1.0/1.0 | 1.0/1.0/1.0 | ||||
Untranslated Text | 0.9745 | 0.987 | 0.625 / 0.5 / 0.8333 | 0.7273 / 1.0 / 0.5714 | ||||
Confusing Answers | 0.966 | 0.9827 | 0.6923 / 0.5625 / 0.9 | 0.7778 / 0.6364 / 1.0 | ||||
Stiffness | 0.9787 | 0.9481 | 0.7368 / 0.6364 / 0.875 | 0.6842 / 1.0 / 0.52 | ||||
Repetitive Expression | 0.9745 | 0.961 | 0.75 / 0.6 / 1.0 | 0.5263 / 0.3571 / 1.0 | ||||
Subject Imprecision | 0.9277 | 0.9394 | 0.7213 / 0.7857 / 0.6667 | 0.72 / 0.8182 / 0.6429 |
Method | task | label | Spearman/Pearson/Kendall | Acc | F1/Precision/Recall | |||
eval | test | eval | test | eval | test | |||
Standard Prompt Paradigm | Sentiment Analysis | Not Meeting the Requirements | 0.9579 0.9564 0.9545 | 0.9693 0.9643 0.9584 | 0.9649 | 0.9279 | 0.875 / 0.875 / 0.875 | 0.7143 / 0.625 / 0.8333 |
Incorrect Answer/Unrelated Matching Results | 0.9825 | 0.9189 | 0.9792 / 1.0 / 0.9592 | 0.9091 / 0.9375 / 0.8824 | ||||
Refusal to Answer | 0.9825 | 0.982 | 0.75/1.0/0.6 | 0.6667 / 0.5 / 1.0 | ||||
Untranslated Text | 1 | 1 | 1.0/1.0/1.0 | 1.0/1.0/1.0 | ||||
Confusing Answers | 0.9912 | 0.982 | 0.8 / 1.0 / 0.6667 | 0.6667 / 0.6667 / 0.6667 | ||||
Stiffness | 0.9825 | 0.964 | 0.875 / 1.0 / 0.7778 | 0.8/0.6667/1.0 | ||||
Repetitive Expression | 0.9912 | 0.9369 | 0.9412 / 1.0 / 0.8889 | 0.6667 / 0.5385 / 0.875 | ||||
Subject Imprecision | 0.9123 | 0.8829 | 0.7727 / 0.9444 / 0.6538 | 0.48 / 0.3333 / 0.8571 | ||||
Knowledge QA | Incorrect Answer/Unrelated Matching Results | 0.4945 0.5153 0.4734 | 0.5063 0.4982 0.48 | 0.8298 | 0.7737 | 0.52 / 0.3611 / 0.9286 | 0.2439 / 0.1667 / 0.4545 | |
Refusal to Answer | 1 | 1 | 1.0/1.0/1.0 | 1.0/1.0/1.0 | ||||
Untranslated Text | 0.9929 | 0.9927 | 0.8889/0.8/1.0 | 0.8571 / 1.0 / 0.75 | ||||
Confusing Answers | 0.9787 | 1 | 0.6667 / 0.5 / 1.0 | 1.0/1.0/1.0 | ||||
Incomplete Answers | 0.7589 | 0.7372 | 0.575 / 0.8846 / 0.4259 | 0.625 / 0.9677 / 0.4615 | ||||
Stiffness | 0.9645 | 1 | 0.6154 / 0.8 / 0.5 | 1.0/1.0/1.0 | ||||
Repetitive Expression | 0.9929 | 0.9781 | 0.9333 / 0.875 / 1.0 | 0.7692 / 0.7143 / 0.8333 | ||||
Subject Imprecision | 0.9574 | 0.9562 | 0.7273 / 0.8 / 0.6667 | 0.5/0.6/0.4286 | ||||
Search QA | Incorrect Answer/Unrelated Matching Results | 0.8263 0.8223 0.7951 | 0.8487 0.8487 0.8352 | 0.9528 | 0.9421 | 0.7692 / 0.9091 / 0.6667 | 0.7586 / 0.9167 / 0.6471 | |
External Links or Diversions | 1 | 0.9835 | 1.0/1.0/1.0 | 0.8333 / 0.7143 / 1.0 | ||||
Refusal to Answer | 0.9685 | 1 | 0.7143 / 0.5556 / 1.0 | 1.0/1.0/1.0 | ||||
Incomplete Answers | 0.8819 | 0.9091 | 0.4444 / 0.4615 / 0.4286 | 0.4211 / 1.0 / 0.2667 | ||||
Stiffness | 0.9606 | 0.9835 | 0.4444 / 0.2857 / 1.0 | 0.75/1.0/0.6 | ||||
Repetitive Expression | 0.9843 | 0.9917 | 0.6667 / 0.6667 / 0.6667 | 0.8/0.6667/1.0 | ||||
Subject Imprecision | 0.9449 | 0.9256 | 0.3636 / 0.2857 / 0.5 | 0.4706/0.8/0.3333 | ||||
Title Generation | Not Meeting the Requirements | 0.9207 0.9322 0.915 | 0.9006 0.9055 0.8963 | 0.9574 | 0.9524 | 0.9695 / 0.9636 / 0.9755 | 0.9657 / 0.9873 / 0.9451 | |
Incorrect Answer/Unrelated Matching Results | 0.9319 | 0.8788 | 0.8298 / 0.907 / 0.7647 | 0.7544 / 0.8958 / 0.6515 | ||||
External Links or Diversion | 1 | 1 | 1.0 / 1.0 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Untranslated Text | 0.9745 | 1 | 0.6667 / 0.5 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Confusing Answers | 0.983 | 0.9913 | 0.8 / 0.8 / 0.8 | 0.8571 / 0.8571 / 0.8571 | ||||
Stiffness | 0.9404 | 0.9827 | 0.5333 / 0.3636 / 1.0 | 0.913 / 1.0 / 0.84 | ||||
Repetitive Expression | 0.9745 | 0.9913 | 0.6667 / 0.6667 / 0.6667 | 0.8333 / 0.7143 / 1.0 | ||||
Subject Imprecision | 0.9277 | 0.9567 | 0.7213 / 0.7857 / 0.6667 | 0.7917/0.95/0.6786 |
Method | task | label | Spearman/Pearson/Kendall | Acc | F1/Precision/Recall | |||
eval | test | eval | test | eval | test | |||
Multi-turn with Zero-shot | Sentiment Analysis | Not Meeting the Requirements | 0.9536 0.9506 0.9467 | 0.9485 0.943 0.9378 | 0.9825 | 0.8919 | 0.9412 / 0.8889 / 1.0 | 0.6471 / 0.5 / 0.9167 |
Incorrect Answer/Unrelated Matching Results | 0.9825 | 0.9369 | 0.9792 / 1.0 / 0.9592 | 0.932 / 0.9231 / 0.9412 | ||||
Refusal to Answer | 0.9737 | 0.9459 | 0.7692 / 0.625 / 1.0 | 0.4 / 0.25 / 1.0 | ||||
Untranslated Text | 0.9912 | 0.991 | 0.9333 / 0.875 / 1.0 | 0.9231 / 1.0 / 0.8571 | ||||
Confusing Answers | 0.9912 | 0.991 | 0.8/1.0/0.6667 | 0.8 / 1.0 / 0.6667 | ||||
Stiffness | 0.9825 | 0.964 | 0.9 / 0.8182 / 1.0 | 0.8/0.6667/1.0 | ||||
Repetitive Expression | 0.9912 | 0.955 | 0.9412 / 1.0 / 0.8889 | 0.7368 / 0.6364 / 0.875 | ||||
Subject Imprecision | 0.9825 | 0.8649 | 0.963 / 0.9286 / 1.0 | 0.4444 / 0.3 / 0.8571 | ||||
Knowledge QA | Incorrect Answer/Unrelated Matching Results | 0.8597 0.8542 0.8487 | 0.7089 0.6794 0.6898 | 0.9645 | 0.9051 | 0.8148 / 0.8462 / 0.7857 | 0.5185 / 0.4375 / 0.6364 | |
Refusal to Answer | 0.9929 | 1 | 0.8/1.0/0.6667 | 1.0/1.0/1.0 | ||||
Untranslated Text | 0.9929 | 1 | 0.8889/0.8/1.0 | 1.0/1.0/1.0 | ||||
Confusing Answers | 0.9929 | 0.9927 | 0.8 / 1.0 / 0.6667 | 0.8889/1.0/0.8 | ||||
Incomplete Answers | 0.9433 | 0.9197 | 0.9245 / 0.9423 / 0.9074 | 0.912 / 0.95 / 0.8769 | ||||
Stiffness | 0.9787 | 0.9854 | 0.8 / 0.8571 / 0.75 | 0.875 / 0.7778 / 1.0 | ||||
Repetitive Expression | 1 | 0.9854 | 1.0/1.0/1.0 | 0.8333 / 0.8333 / 0.8333 | ||||
Subject Imprecision | 0.9716 | 0.9635 | 0.8182 / 0.9 / 0.75 | 0.6154 / 0.6667 / 0.5714 | ||||
Search QA | Incorrect Answer/Unrelated Matching Results | 0.8574 0.8566 0.8377 | 0.9438 0.9434 0.9394 | 0.9449 | 0.9835 | 0.7586 / 0.7857 / 0.7333 | 0.9412 / 0.9412 / 0.9412 | |
External Links or Diversions | 1 | 1 | 1.0 / 1.0 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Refusal to Answer | 0.9843 | 1 | 0.8333 / 0.7143 / 1.0 | 1.0/1.0/1.0 | ||||
Incomplete Answers | 0.9213 | 0.9917 | 0.7222 / 0.5909 / 0.9286 | 0.9677 / 0.9375 / 1.0 | ||||
Stiffness | 0.9685 | 0.9917 | 0.5/0.3333/1.0 | 0.8889/1.0/0.8 | ||||
Repetitive Expression | 0.9843 | 0.9917 | 0.6667 / 0.6667 / 0.6667 | 0.6667 / 1.0 / 0.5 | ||||
Subject Imprecision | 0.9528 | 0.9835 | 0.5714 / 0.4 / 1.0 | 0.9231 / 0.8571 / 1.0 | ||||
Title Generation | Not Meeting the Requirements | 0.915 0.9142 0.9124 | 0.9206 0.9243 0.9148 | 0.983 | 0.9697 | 0.9878 / 0.9818 / 0.9939 | 0.9785 / 0.9876 / 0.9695 | |
Incorrect Answer/Unrelated Matching Results | 0.8681 | 0.9957 | 0.7597 / 0.6282 / 0.9608 | 0.9925 / 0.9851 / 1.0 | ||||
External Links or Diversion | 1 | 1 | 1.0 / 1.0 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Untranslated Text | 0.983 | 1 | 0.75 / 0.6 / 1.0 | 1.0/1.0/1.0 | ||||
Confusing Answers | 1 | 1 | 1.0 / 1.0 / 1.0 | 1.0 / 1.0 / 1.0 | ||||
Stiffness | 0.9872 | 1 | 0.8421 / 0.7273 / 1.0 | 1.0/1.0/1.0 | ||||
Repetitive Expression | 0.9957 | 1 | 0.9474 / 0.9 / 1.0 | 1.0/1.0/1.0 | ||||
Subject Imprecision | 0.9489 | 0.9957 | 0.8462 / 0.7333 / 1.0 | 0.9825 / 0.9655 / 1.0 |
Method | task | label | Spearman/Pearson/Kendall | Acc | F1/Precision/Recall | |||
eval | test | eval | test | eval | test | |||
Qwen-72B-Chat + SFT | Sentiment Analysis | Not Meeting the Requirements | 0.4565 0.4166 0.4366 | 0.4211 0.4308 0.4025 | 0.8158 | 0.7297 | 0.16 / 0.2222 / 0.125 | 0.2105 / 0.1538 / 0.3333 |
Incorrect Answer/Unrelated Matching Results | 0.8158 | 0.8108 | 0.7586 / 0.8684 / 0.6735 | 0.764 / 0.8947 / 0.6667 | ||||
Refusal to Answer | 0.9386 | 0.9459 | 0.5333 / 0.4 / 0.8 | 0.4 / 0.25 / 1.0 | ||||
Untranslated Text | 0.8772 | 0.7928 | 0.4167 / 0.2941 / 0.7143 | 0.303 / 0.1923 / 0.7143 | ||||
Confusing Answers | 0.8596 | 0.8468 | 0.2 / 0.1176 / 0.6667 | 0.1905 / 0.1111 / 0.6667 | ||||
Stiffness | 0.5702 | 0.4865 | 0.1967 / 0.1154 / 0.6667 | 0.1739 / 0.0984 / 0.75 | ||||
Repetitive Expression | 0.4737 | 0.4865 | 0.1892 / 0.1077 / 0.7778 | 0.1972 / 0.1111 / 0.875 | ||||
Subject Imprecision | 0.5877 | 0.6216 | 0.3188 / 0.2558 / 0.4231 | 0.087 / 0.0513 / 0.2857 | ||||
Knowledge QA | Incorrect Answer/Unrelated Matching Results | 0.1886 0.1896 0.1742 | 0.2444 0.2142 0.2301 | 0.7305 | 0.5766 | 0.3214 / 0.2143 / 0.6429 | 0.1471 / 0.0877 / 0.4545 | |
Refusal to Answer | 0.922 | 0.927 | 0.3529 / 0.2143 / 1.0 | 0.375 / 0.2308 / 1.0 | ||||
Untranslated Text | 0.7305 | 0.7737 | 0.0952 / 0.0526 / 0.5 | 0.1143 / 0.0645 / 0.5 | ||||
Confusing Answers | 0.8865 | 0.8248 | 0.2 / 0.1176 / 0.6667 | 0.2 / 0.12 / 0.6 | ||||
Incomplete Answers | 0.5887 | 0.5985 | 0.4528 / 0.4615 / 0.4444 | 0.56 / 0.5833 / 0.5385 | ||||
Stiffness | 0.617 | 0.6058 | 0.1562 / 0.0893 / 0.625 | 0.1818 / 0.1017 / 0.8571 | ||||
Repetitive Expression | 0.4468 | 0.438 | 0.1333 / 0.0723 / 0.8571 | 0.1348 / 0.0723 / 1.0 | ||||
Subject Imprecision | 0.6454 | 0.6423 | 0.2647 / 0.1607 / 0.75 | 0.1695 / 0.0962 / 0.7143 | ||||
Search QA | Incorrect Answer/Unrelated Matching Results | 0.5705 0.5398 0.5226 | 0.4772 0.4553 0.4425 | 0.748 | 0.7934 | 0.2381 / 0.1852 / 0.3333 | 0.4186 / 0.3462 / 0.5294 | |
External Links or Diversions | 0.8425 | 0.8678 | 0.3333 / 0.2273 / 0.625 | 0.2 / 0.1333 / 0.4 | ||||
[…]

### v2

@@ Appendix B Experimental Results
The details of all experimental results are tabulated in Table 10 to Table 16.
Method
task
label
Spearman/Pearson/Kendall
Acc
F1/Precision/Recall
eval
test
eval
test
eval
test
Arbitrariness
Sentiment Analysis
Not Meeting the Requirements
0.9551
0.9523
0.9484
0.9515
0.943
0.9268
0.9123
0.8829
0.75 / 0.625 / 0.9375
0.6286 / 0.4783 / 0.9167
Incorrect Answer/Unrelated Matching Results
0.9737
0.9279
0.9684 / 1.0 / 0.9388
0.92 / 0.9388 / 0.902
Refusal to Answer
0.9737
0.964
0.6667 / 0.75 / 0.6
0.5 / 0.3333 / 1.0
Untranslated Text
0.9825
0.982
0.875 / 0.7778 / 1.0
0.875 / 0.7778 / 1.0
Confusing Answers
0.9912
0.982
0.8571 / 0.75 / 1.0
0.75 / 0.6 / 1.0
Stiffness
0.9649
0.9279
0.8182 / 0.6923 / 1.0
0.6667 / 0.5 / 1.0
Repetitive Expression
0.9649
0.955
0.7143 / 1.0 / 0.5556
0.5455 / 1.0 / 0.375
Subject Imprecision
0.8947
0.7838
0.7857 / 0.7333 / 0.8462
0.3333 / 0.2069 / 0.8571
Knowledge QA
Incorrect Answer/Unrelated Matching Results
0.5242
0.5345
0.4974
0.5459
0.5436
0.518
0.8085
0.7591
0.4906 / 0.3333 / 0.9286
0.2979 / 0.1944 / 0.6364
Refusal to Answer
1
0.9927
1.0/1.0/1.0
0.8571 / 0.75 / 1.0
Untranslated Text
0.9858
0.9927
0.75 / 0.75 / 0.75
0.8571 / 1.0 / 0.75
Confusing Answers
0.9858
1
0.75 / 0.6 / 1.0
1.0/1.0/1.0
Incomplete Answers
0.7589
0.6934
0.6136 / 0.7941 / 0.5
0.5625 / 0.871 / 0.4154
Stiffness
0.9716
1
0.7143 / 0.8333 / 0.625
1.0/1.0/1.0
Repetitive Expression
0.9858
0.9781
0.8571 / 0.8571 / 0.8571
0.7273 / 0.8 / 0.6667
Subject Imprecision
0.9645
0.9416
0.7619 / 0.8889 / 0.6667
0.3333 / 0.4 / 0.2857
Search QA
Incorrect Answer/Unrelated Matching Results
0.7946
0.7901
0.7852
0.814
0.818
0.791
0.9449
0.9587
0.7586 / 0.7857 / 0.7333
0.8387 / 0.9286 / 0.7647
External Links or Diversions
0.9921
0.9752
0.9333 / 1.0 / 0.875
0.7273 / 0.6667 / 0.8
Refusal to Answer
0.9685
1
0.7143 / 0.5556 / 1.0
1.0 / 1.0 / 1.0
Incomplete Answers
0.8661
0.9174
0.32 / 0.3636 / 0.2857
0.5455 / 0.8571 / 0.4
Stiffness
0.9606
0.9752
0.4444 / 0.2857 / 1.0
0.6667 / 0.75 / 0.6
Repetitive Expression
0.9528
0.9421
0.4 / 0.2857 / 0.6667
0.3636 / 0.2222 / 1.0
Subject Imprecision
0.9528
0.9174
0.5714 / 0.4 / 1.0
0.5455 / 0.6 / 0.5
Title Generation
Not Meeting the Requirements
0.535
0.5629
0.5153
0.5729
0.593
0.5483
0.617
0.6147
0.6457 / 0.7387 / 0.5734
0.6426 / 0.7477 / 0.5634
Incorrect Answer/Unrelated Matching Results
0.9064
0.8745
0.8036 / 0.7377 / 0.8824
0.7752 / 0.7937 / 0.7576
External Links or Diversion
1
1
1.0/1.0/1.0
1.0/1.0/1.0
Untranslated Text
0.9745
0.987
0.625 / 0.5 / 0.8333
0.7273 / 1.0 / 0.5714
Confusing Answers
0.966
0.9827
0.6923 / 0.5625 / 0.9
0.7778 / 0.6364 / 1.0
Stiffness
0.9787
0.9481
0.7368 / 0.6364 / 0.875
0.6842 / 1.0 / 0.52
Repetitive Expression
0.9745
0.961
0.75 / 0.6 / 1.0
0.5263 / 0.3571 / 1.0
Subject Imprecision
0.9277
0.9394
0.7213 / 0.7857 / 0.6667
0.72 / 0.8182 / 0.6429
Table 10: The overall results of Arbitrariness.
Method
task
label
Spearman/Pearson/Kendall
Acc
F1/Precision/Recall
eval
test
eval
test
eval
test
Standard PromptParadigm
Sentiment Analysis
Not Meeting the Requirements
0.9579
0.9564
0.9545
0.9693
0.9643
0.9584
0.9649
0.9279
0.875 / 0.875 / 0.875
0.7143 / 0.625 / 0.8333
Incorrect Answer/Unrelated Matching Results
0.9825
0.9189
0.9792 / 1.0 / 0.9592
0.9091 / 0.9375 / 0.8824
Refusal to Answer
0.9825
0.982
0.75/1.0/0.6
0.6667 / 0.5 / 1.0
Untranslated Text
1
1
1.0/1.0/1.0
1.0/1.0/1.0
Confusing Answers
0.9912
0.982
0.8 / 1.0 / 0.6667
0.6667 / 0.6667 / 0.6667
Stiffness
0.9825
0.964
0.875 / 1.0 / 0.7778
0.8/0.6667/1.0
Repetitive Expression
0.9912
0.9369
0.9412 / 1.0 / 0.8889
0.6667 / 0.5385 / 0.875
Subject Imprecision
0.9123
0.8829
0.7727 / 0.9444 / 0.6538
0.48 / 0.3333 / 0.8571
Knowledge QA
Incorrect Answer/Unrelated Matching Results
0.4945
0.5153
0.4734
0.5063
0.4982
0.48
0.8298
0.7737
0.52 / 0.3611 / 0.9286
0.2439 / 0.1667 / 0.4545
Refusal to Answer
1
1
1.0/1.0/1.0
1.0/1.0/1.0
Untranslated Text
0.9929
0.9927
0.8889/0.8/1.0
0.8571 / 1.0 / 0.75
Confusing Answers
0.9787
1
0.6667 / 0.5 / 1.0
1.0/1.0/1.0
Incomplete Answers
0.7589
0.7372
0.575 / 0.8846 / 0.4259
0.625 / 0.9677 / 0.4615
Stiffness
0.9645
1
0.6154 / 0.8 / 0.5
1.0/1.0/1.0
Repetitive Expression
0.9929
0.9781
0.9333 / 0.875 / 1.0
0.7692 / 0.7143 / 0.8333
Subject Imprecision
0.9574
0.9562
0.7273 / 0.8 / 0.6667
0.5/0.6/0.4286
Search QA
Incorrect Answer/Unrelated Matching Results
0.8263
0.8223
0.7951
0.8487
0.8487
0.8352
0.9528
0.9421
0.7692 / 0.9091 / 0.6667
0.7586 / 0.9167 / 0.6471
External Links or Diversions
1
0.9835
1.0/1.0/1.0
0.8333 / 0.7143 / 1.0
Refusal to Answer
0.9685
1
0.7143 / 0.5556 / 1.0
1.0/1.0/1.0
Incomplete Answers
0.8819
0.9091
0.4444 / 0.4615 / 0.4286
0.4211 / 1.0 / 0.2667
Stiffness
0.9606
0.9835
0.4444 / 0.2857 / 1.0
0.75/1.0/0.6
Repetitive Expression
0.9843
0.9917
0.6667 / 0.6667 / 0.6667
0.8/0.6667/1.0
Subject Imprecision
0.9449
0.9256
0.3636 / 0.2857 / 0.5
0.4706/0.8/0.3333
Title Generation
Not Meeting the Requirements
0.9207
0.9322
0.915
0.9006
0.9055
0.8963
0.9574
0.9524
0.9695 / 0.9636 / 0.9755
0.9657 / 0.9873 / 0.9451
Incorrect Answer/Unrelated Matching Results
0.9319
0.8788
0.8298 / 0.907 / 0.7647
0.7544 / 0.8958 / 0.6515
External Links or Diversion
1
1
1.0 / 1.0 / 1.0
1.0 / 1.0 / 1.0
Untranslated Text
0.9745
1
0.6667 / 0.5 / 1.0
1.0 / 1.0 / 1.0
Confusing Answers
0.983
0.9913
0.8 / 0.8 / 0.8
0.8571 / 0.8571 / 0.8571
Stiffness
0.9404
0.9827
0.5333 / 0.3636 / 1.0
0.913 / 1.0 / 0.84
Repetitive Expression
0.9745
0.9913
0.6667 / 0.6667 / 0.6667
0.8333 / 0.7143 / 1.0
Subject Imprecision
0.9277
0.9567
0.7213 / 0.7857 / 0.6667
0.7917/0.95/0.6786
Table 11: The overall results of Standard Prompt Paradigm
Method
task
label
Spearman/Pearson/Kendall
Acc
F1/Precision/Recall
eval
test
eval
test
eval
test
Multi-turnwith Zero-shot
Sentiment Analysis
Not Meeting the Requirements
0.9536
0.9506
0.9467
0.9485
0.943
0.9378
0.9825
0.8919
0.9412 / 0.8889 / 1.0
0.6471 / 0.5 / 0.9167
Incorrect Answer/Unrelated Matching Results
0.9825
0.9369
0.9792 / 1.0 / 0.9592
0.932 / 0.9231 / 0.9412
Refusal to Answer
0.9737
0.9459
0.7692 / 0.625 / 1.0
0.4 / 0.25 / 1.0
Untranslated Text
0.9912
0.991
0.9333 / 0.875 / 1.0
0.9231 / 1.0 / 0.8571
Confusing Answers
0.9912
0.991
0.8/1.0/0.6667
0.8 / 1.0 / 0.6667
Stiffness
0.9825
0.964
0.9 / 0.8182 / 1.0
0.8/0.6667/1.0
Repetitive Expression
0.9912
0.955
0.9412 / 1.0 / 0.8889
0.7368 / 0.6364 / 0.875
Subject Imprecision
0.9825
0.8649
0.963 / 0.9286 / 1.0
0.4444 / 0.3 / 0.8571
Knowledge QA
Incorrect Answer/Unrelated Matching Results
0.8597
0.8542
0.8487
0.7089
0.6794
0.6898
0.9645
0.9051
0.8148 / 0.8462 / 0.7857
0.5185 / 0.4375 / 0.6364
Refusal to Answer
0.9929
1
0.8/1.0/0.6667
1.0/1.0/1.0
Untranslated Text
0.9929
1
0.8889/0.8/1.0
1.0/1.0/1.0
Confusing Answers
0.9929
0.9927
0.8 / 1.0 / 0.6667
0.8889/1.0/0.8
Incomplete Answers
0.9433
0.9197
0.9245 / 0.9423 / 0.9074
0.912 / 0.95 / 0.8769
Stiffness
0.9787
0.9854
0.8 / 0.8571 / 0.75
0.875 / 0.7778 / 1.0
Repetitive Expression
1
0.9854
1.0/1.0/1.0
0.8333 / 0.8333 / 0.8333
Subject Imprecision
0.9716
0.9635
0.8182 / 0.9 / 0.75
0.6154 / 0.6667 / 0.5714
Search QA
Incorrect Answer/Unrelated Matching Results
0.8574
0.8566
0.8377
0.9438
0.9434
0.9394
0.9449
0.9835
0.7586 / 0.7857 / 0.7333
0.9412 / 0.9412 / 0.9412
External Links or Diversions
1
1
1.0 / 1.0 / 1.0
1.0 / 1.0 / 1.0
Refusal to Answer
0.9843
1
0.8333 / 0.7143 / 1.0
1.0/1.0/1.0
Incomplete Answers
0.9213
0.9917
0.7222 / 0.5909 / 0.9286
0.9677 / 0.9375 / 1.0
Stiffness
0.9685
0.9917
0.5/0.3333/1.0
0.8889/1.0/0.8
Repetitive Expression
0.9843
0.9917
0.6667 / 0.6667 / 0.6667
0.6667 / 1.0 / 0.5
Subject Imprecision
0.9528
0.9835
0.5714 / 0.4 / 1.0
0.9231 / 0.8571 / 1.0
Title Generation
Not Meeting the Requirements
0.915
0.9142
0.9124
0.9206
0.9243
0.9148
0.983
0.9697
0.9878 / 0.9818 / 0.9939
0.9785 / 0.9876 / 0.9695
Incorrect Answer/Unrelated Matching Results
0.8681
0.9957
0.7597 / 0.6282 / 0.9608
0.9925 / 0.9851 / 1.0
External Links or Diversion
1
1
1.0 / 1.0 / 1.0
1.0 / 1.0 / 1.0
Untranslated Text
0.983
1
0.75 / 0.6 / 1.0
1.0/1.0/1.0
Confusing Answers
1
1
1.0 / 1.0 / 1.0
1.0 / 1.0 / 1.0
Stiffness
0.9872
1
0.8421 / 0.7273 / 1.0
1.0/1.0/1.0
Repetitive Expression
0.9957
1
0.9474 / 0.9 / 1.0
1.0/1.0/1.0
Subject Imprecision
0.9489
0.9957
0.8462 / 0.7333 / 1.0
0.9825 / 0.9655 / 1.0
Table 12: The overall results of Multi-turn with Zero-shot
Method
task
label
Spearman/Pearson/Kendall
Acc
F1/Precision/Recall
eval
test
eval
test
eval
test
Qwen-72B-Chat+ SFT
Sentiment Analysis
Not Meeting the Requirements
0.4565
0.4166
0.4366
0.4211
0.4308
0.4025
0.8158
0.7297
0.16 / 0.2222 / 0.125
0.2105 / 0.1538 / 0.3333
Incorrect Answer/Unrelated Matching Results
0.8158
0.8108
0.7586 / 0.8684 / 0.6735
0.764 / 0.8947 / 0.6667
Refusal to Answer
0.9386
0.9459
0.5333 / 0.4 / 0.8
0.4 / 0.25 / 1.0
Untranslated Text
0.8772
0.7928
0.4167 / 0.2941 / 0.7143
0.303 / 0.1923 / 0.7143
Confusing Answers
0.8596
0.8468
0.2 / 0.1176 / 0.6667
0.1905 / 0.1111 / 0.6667
Stiffness
0.5702
0.4865
0.1967 / 0.1154 / 0.6667
0.1739 / 0.0984 / 0.75
Repetitive Expression
0.4737
0.4865
0.1892 / 0.1077 / 0.7778
0.1972 / 0.1111 / 0.875
Subject Imprecision
0.5877
0.6216
0.3188 / 0.2558 / 0.4231
0.087 / 0.0513 / 0.2857
Knowledge QA
Incorrect Answer/Unrelated Matching Results
0.1886
0.1896
0.1742
0.2444
0.2142
0.2301
0.7305
0.5766
0.3214 / 0.2143 / 0.6429
0.1471 / 0.0877 / 0.4545
Refusal to Answer
0.922
0.927
0.3529 / 0.2143 / 1.0
0.375 / 0.2308 / 1.0
Untranslated Text
0.7305
0.7737
0.0952 / 0.0526 / 0.5
0.1143 / 0.0645 / 0.5
Confusing Answers
0.8865
0.8248
0.2 / 0.1176 / 0.6667
0.2 / 0.12 / 0.6
Incomplete Answers
0.5887
0.5985
0.4528 / 0.4615 / 0.4444
0.56 / 0.5833 / 0.5385
Stiffness
0.617
0.6058
0.1562 / 0.0893 / 0.625
0.1818 / 0.1017 / 0.8571
Repetitive Expression
0.4468
0.438
0.1333 / 0.0723 / 0.8571
0.1348 / 0.0723 / 1.0
Subject Imprecision
0.6454
0.6423
0.2647 / 0.1607 / 0.75
0.1695 / 0.0962 / 0.7143
Search QA
Incorrect Answer/Unrelated Matching Results
0.5705
0.5398
0.5226
0.4772
0.4553
0.4425
0.748
0.7934
0.2381 / 0.1852 / 0.3333
0.4186 / 0.3462 / 0.5294
External Links or Diversions
0.8425
0.8678
0.3333 / 0.2273 / 0.625
0.2 / 0.1333 / 0.4
Refusal to Answer
0.9213
0.9504
[…]